A jar contains 3 red, 5 green, 2 blue and 6 yellow marbles. A marble is chosen at random from the jar. After replacing it, a second marble is chosen. What is the probability of choosing a green and then a yellow marble?


1
Expert's answer
2021-06-14T19:16:09-0400
"3+5+2+6=16"


"P(first\ G, second\ Y)=\dfrac{5}{16}\cdot\dfrac{6}{16}=\dfrac{15}{128}"



The probability of choosing a green and then a yellow marble is "\dfrac{15}{128}."